### v1

@@ Abstract
Describing Li-ion battery cathodes in terms of distinct transition metal or oxygen redox regimes can lead to confusion in understanding metal-ligand hybridisation, oxygen dimerisation, and degradation. There is a pressing need to study the electronic structure of these materials and determine the role each cation and anion plays in charge compensation. Here, we employ transition metal L-edge X-ray Resonance Photoemission Spectroscopy in conjunction with Single Impurity Anderson simulations to directly evaluate the redox mechanisms in (de-)lithiated battery electrodes. This approach reconciles the redox description of two canonical cathodes—LiMn0.6Fe0.4PO4 and LiNiO2—in terms of varying degrees of charge transfer using the established Zaanen-Sawatzky-Allen framework, common to condensed matter physics. In LiMn0.6Fe0.4PO4, we show that capacity arises due to the depopulation of metal 3d states. Whereas, in LiNiO2, charge transfer dominates and redox occurs through the formation and elimination of ligand hole states. This work clarifies the role of oxygen in Ni-rich system and provides a framework to explain how capacity can be extracted from oxygen-dominated states in highly covalent systems without needing to invoke dimerisation.
@@
Li-ion battery cathodes, Electronic structure, X-ray spectroscopy, X-ray simulationsState-of-the-art Ni-rich cathodes are currently voltage limited in order to suppress oxygen loss and increase cycling lifetimes. [1, 2] Understanding oxygen loss can significantly improve Ni-rich cathode performance. Oxygen loss and oxygen redox (involving O2 formation) are often conflated, [3, 4] and points to the need for a better understanding of oxygen participation more generally. Over the last decade, oxygen redox conceptually distinguishes between non-bonding O 2p orbitals and metal-ligand orbital hybridisation. [5] This has been solidified by studies of Li-excess cathodes where oxygen oxidation is described in terms of O2 formation from resonant inelastic x-ray scattering (RIXS) experiments.[4] Stoichiometric layered transition metal (TM) oxide systems do not deliver excess capacity beyond that which can be explained by metal redox alone. However, features associated with oxygen dimers using RIXS have been reported at the top of charge for multiple layered oxides. [6, 7, 8, 9] More recent RIXS studies by Tarascon et al. have raised doubts on whether RIXS signals correlate to reversible electrochemical capacity in any of the systems which exhibit these O2 dimer features.[10] Nevertheless, beyond density functional theory studies have now confirmed that oxygen does participate via TM-oxygen hybridisation right from the onset of delithiation, [11] consistent with a metal-ligand redox mechanism that can be used to contextualise operando X-ray studies.[12] Moreover, structure disorder rather than O2 formation has been proposed to play a more prominent role in defining the electrochemistry of layered materials. [13, 14] As a result, further investigation of oxygen charge compensation in Ni-rich materials via a direct characterisation of the charge-compensating orbitals during intercalation reactions is needed.
Equation 1 illustrates the end points of delithiation reactions of LiNiO2 and its concurrent charge compensation that leads to a higher Ni oxidation state. What this Ni oxidation state evolution means as a function of the LiNiO2 charge compensation mechanism can guide us to identify the root cause of O-loss. The classical redox model ascribes oxidation states to the charge density of TM 3d orbitals based on the assumption of a fully ionic character of the TM-O bonds, [15, 16] leading to the conventional view of oxidation state evolution conceived as discrete changes in the occupancy of the TM 3d orbitals. [17] Yet increasing experimental evidence invites us to re-evaluate the suitability of this ionic picture for LiNiO2 since it has been shown its electronic structure can be strongly influenced by complex factors beyond the ionic description, including: bond-disproportionate sites, [18] intersite charge fluctuations, [19] and metal-ligand orbital hybridisation. [20] Seeking an understanding of redox mechanisms where these more complex electronic features are better accounted for has been discussed in theoretical contexts [21, 14, 22] but requires further experimental validation to unveil how these factors can determine the redox nature and the mechanism for oxygen participation during charge compensation in Ni-rich layered materials.
Most studies in Ni-rich layered compounds to date focus on indirect experimental probe of the charge compensation mechanism. The mechanism is inferred by probing the electronic structure of cathode material on core and/or empty orbitals different than the charge-compensating orbital. In the classical redox view, this indirect approach suffices. However, the complex LiNiO2 electronic features, which can also modulate the character of charge-compensating orbital and potentially the redox mechanism, demand re-interpreting the indirect approach data. For instance, during Li deintercalation of LiNiO2 cathodes, HAXPES Ni 2p peaks narrow their peak width rather than shift in energy. [23] The XANES Ni edge evolution is not linear as expected for a fully ionic character of the Ni-O bonds. [24] Raman Eg and A1g peak intensities increase suggesting a dynamical evolution of covalency of the Ni-O bonds in contrast to the ionic assumption. [19] The energy loss spectrum of Ni L-edge RIXS originates from a Ni 3d8 configuration as opposed to the Ni 3d7 orbital occupancy proposed by the classical redox mechanism for LiNiO2. [19] Cutting-edge computational approaches [25] suggest that in these complex electronic conditions, an experimental method capable to directly probe the charge-compensating orbital character while distinguishing its dependence on bond-disproportionate sites, intersite charge fluctuation, and metal-ligand orbital hybridisation represents an opportunity for a substantial leap towards unveiling the nature of oxygen redox.
We close this gap by implementing TM L-edge X-ray Absorption Spectroscopy (XAS) assisted Resonant Photoemission Spectroscopy (RPES). [26] In typical valence-band photoemission (VB-PES) study, we fix the detection window in the range dominated by the charge-compensating orbitals. However, in this setting, the spectra do not contain element-specific information. TM L-edge XAS-assisted RPES, i.e. VB-PES with a tuned X-ray beam energy to resonate at a TM L-edge XAS transition, [27] not only enables this element sensitivity but filters out spectral signals whose nature is not the electronic feature at the resonating TM L-edge XAS transition. Single impurity Anderson (SIA) simulations [28] of TM L-edge XAS can identify the nature of the TM L-edge XAS peaks, which can be of the t2g and eg orbital, intersite charge fluctuation, bond-disproportionate site, and/or metal-ligand orbital hybridisation kind. Thus, by selectively resonating at some of these peaks, we highlight, onto the charge-compensating orbital spectrum, spectral weight attributed to those factors and, thereby, determine to what extent they influence the oxygen redox mechanism at different states of charges (SoCs).
@@
Here, we used the TM L-edge XAS-assisted RPES approach to directly study charge compensation mechanisms in the context of the simple ionic picture and more complex electronic structures. We used this approach on two canonical systems, LiMn0.6Fe0.4PO4 (LMFP64) and LiNiO2 (LNO). The high quality and electrochemical performance of the LMFP64 material and electrodes used in this work have been proven previously. [29] The LNO electrodes of this work exhibit the expected accelerated electrochemical performance decay [30] as a consequence of cycling above its O-loss voltage threshold at 4.1 V vs. Li+/Li, which we used to intentionally condition the material with relevant voltage windows for this work. LMFP64 shows the classical redox picture due to an intense induction effect from the P ions. [31] In contrast, by carefully considering bond-disproportionate site, intersite charge fluctuation, and metal-ligand orbital hybridisation effects, a different redox view emerges in LNO more linked to a local site and ligand hole reconfiguration. Our findings provide critical insights to the electronic structure and a pathway to understanding oxygen redox.
The Classical Redox in Phosphate Materials
@@
Similar to equation 1, equation 2 shows the oxidation state evolution of phosphate cathode materials as a result of delithiation. However, contrary to the LNO case, this formal oxidation state evolution largely reflects electron counting of the Fe 3d-orbital because of LMPF64 electronic structure is within the ionic approximation. Namely, the Fe2 oxidation state translates to six electrons on the Fe 3d-orbital and during delithiation the electron occupancy must decrease by one unit for the Fe ion to evolve from the formal 2+ to the 3+ oxidation state. To validate this view under our approach, we prepared a set of LMFP64 half-coin cell formats to charge the cells to targeted SoC as indicated by the coloured circles in Figure 1a. In general, LMFP shows unexpected structural transformations with fast charge rates. [32] Those transformations can influence the intercalation reactions and consequently the rate at which Fe ions undergo charge compensation. [33] To avoid those and have a clear cut of the redox mechanism, the slow C/20 charging rate was used. This guarantees a voltage profile with negligible polarisation and, as a result, a better approximation to the open circuit voltage profile (the true voltage profile of the material). [3]
@@
The top panel in Figure 1b shows the experimental Fe L-edge spectrum of the pristine electrode (no cycling and no exposure to electrolyte) and the SIA simulated Fe L-edge spectrum. We performed the Fe L-edge simulations without explicit Fe 3d - O ligand 2p hybridisation (to guarantee an ionic picture) and with six Fe 3d electrons (to generate the expected 2+ valence). The remarkable match between the Fe L-edge experimental and simulated spectra confirms the ionic picture best describes the LMFP64 electronic structure; this means that an increase in the oxidation state inherently implies a reduction in the electron occupancy of the Fe 3d orbital. To simulate the 100% SoC condition, we deliberately reduced the number of Fe 3d electrons from six to five in the simulation while keeping the no Fe 3d - O ligand 2p hybridisation condition. Spectral comparison shows again a remarkable match of the simulated and experimental Fe L-edge spectra (bottom panel in Figure 1b). Combining these findings for the pristine and 100% SoC electrodes strongly suggests that the Fe redox mechanism involves a reduction in the number of Fe 3d electrons, leading to an increase in the Fe oxidation state from 2+ to 3+ in LMFP64 during charge. Notice that in the 704 - 709 eV energy range, Fe 3d-orbitals are the dominant feature of the Fe L-edge XAS transitions during charge (Figure 1c). Resonating in this range with RPES experiments will reveal the LMFP64 redox mechanism as a depopulation of the Fe 3d orbitals while the cell charges, proving the classical redox view in LMFP64.
@@
Figure 1d shows RPES results at progressively increasing SoC (pristine or 0, 20, 40, 70 and 100, respectively) in the Fe 3d dominated energy range. With the increasing SoC, we observe the RPES Fe 3d intensity at 2 eV decreases after the Fe2+/3+ potential plateaus in Figure 1a, experimentally confirming the conventional redox picture and that oxidation state changes in LMFP64 links to the valence-counting formalism on the Fe 3d orbital. Notice how for 40% SoC, our results show a mixture of Fe2+ and Fe3+ in both the Fe L-edge and RPES while bulk-sensitive Fe K-edge together with Mössbauer spectra [34] report full transition to Fe3+ state at 40% SoC. The mixed valence can be attributed to involuntary cell discharge during cell dismount from the cyclers. However, our findings are not affected as conventional charge compensation holds true for the full SoC range. A similar procedure was employed to investigate the participation of the Mn sites in the redox mechanism. The analysis prompted further investigation of a possible degradation mechanism stemming from the Mn2+/3+ redox couple that we will explore in a subsequent more in-depth manuscript regarding LMFP64 degradation.
From the Ionic to the Ligand Hole Picture in Nickel Oxides
@@
Before studying redox reactions in lithium nickel oxides in the next section, dedicated to elucidating the influence of TM 3d - O ligand 2p hybridisation and intersite charge fluctuation effects on redox, we need to show how to identify spectroscopically those features using our approach. To demonstrate this, we prepared a set of four Ni-oxide compounds (LaNiO2, NiO, LiNiO2 and Li2NiO3), where the Ni ion is oxidised in different valence states (i.e. Ni1+, Ni2+, Ni3+ and Ni4+). The Ni-O ligand hybridisation is minimal in LaNiO2 [35] while playing a substantial role in NiO. [36] LiNiO2 is better described with intersite charge fluctuations, explicit TM 3d - O ligand 2p hybridisation, and bond-disproportionate NiO6 sites. [23] Li2NiO3 is the endpoint of Li-excess and its electronic structure is better described by intersite charge fluctuations and explicit TM 3d - O ligand 2p hybridisation. [37] Therefore, our approach applied to the Ni oxide material set will show the spectral difference of these complex features.
Figure 2a shows Ni L-edge XAS spectra of LaNiO2, NiO, LiNiO2, and Li2NiO3 and the corresponding RPES measurements in Figure 2b. In contrast to the Fe transition metal case of LMFP64 cathodes, the Ni 3d bands (arrows in Figure 2b) are deep in the energy range between 8 eV and 10 eV. The energy position of these deep bands indicates that Ni redox, i.e. Ni3+(Ni 3d7) to Ni4+(Ni 3d6), [17, 14, 22, 31] implies a minimal cell voltage of at least 8 V. However, experimental observations of nominal cell voltages of approximately 3.8 V vs. Li+/Li for LiNiO2, [30] and Li2NiO3 [37] discredit this view and suggest that charge compensation must occur in shallower energy bands than the Ni3d bands. The white squares in Figure 2b highlight these shallower energy bands that result from resonance within the Ni L-edge 855 to 865 eV region. This scenario makes us question what the nature of these shallower energy bands is. By inspecting the Ni L-edge transitions within the observed range using SIA Ni L-edge simulations of NiO, we provide insights into the nature of those bands.
Figure 2c shows our calculations of NiO Ni L-edge spectra in the ionic picture and with explicit Ni 3d - O ligand 2p hybridisation and intersite charge fluctuation. The hybridisation is implemented within a second quantization formalism between Ni 3d orbitals and symmetry-adapted combinations of O 2p orbitals, which are constructed as linear combinations of O 2p Wannier-like functions that match the symmetry of the Ni 3d orbitals. [28] A clear advantage of using these Wannier-like ligand orbitals over the simplified ligand-field-theory description is that they retain the local symmetry and spatial distribution of the O 2p states in the solid-state environment. The considered intersite charge fluctuation is charge transfer and refers to the possibility for some of the ligand p electrons to pour into the TM d band, leaving a hole in the ligand orbital denoted by , reaching a more stable electronic configuration. The spontaneity of this transfer is given by the factor , with negative values meaning spontaneous charge migration. NiO Ni L-edge spectral simulations are performed with a positive , ensuring that orbital hybridisation is the only dominant feature in the simulations.
The spectral features between 855 and 865 eV are resolved in simulations when we added explicit hybridisation between the O ligand 2p orbitals and the Ni 3d orbitals. We note that a small charge transfer of (0.3 e-) occurs. The mechanism of this charge transfer is not the spontaneous charge transfer as is positive for NiO. This charge transfer emerges from neighbouring oxygen ions that share some charge as a result of the induced X-ray transition, also referred to as Zhang-Rice bound state. [36] Therefore, RPES resonating at these energies intensifies spectral signal from charge transfer effects due to hybridisation. In systems with negative , charge transfer character becomes more relevant in the Ni L-edge energy range between 855 and 865 eV, offering an interpretation to the RPES heatmap spectral region enclosed by the white squares for LiNiO2 and Li2NiO3. These spectral signatures are from ligand hole states. Therefore, the LiNiO2 electronic structure has a strong and stable ligand hole character. LiNiO2 redox is thus excepted to occur on this ligand hole orbitals.
XAS at the O K-edge experimentally confirms the stabilisation of these ligand hole orbitals in LiNiO2 and Li2NiO3. Figure 2d describes the absorption intensity at the energy reference, which was determined by the position of the main peak in the O 1s XPS region, as reported in previous literature. [38] The Ligand Hole region (O K-edge pre-edge) highlighted in Figure 2d displays an increased absorption intensity between NiO and LiNiO2. This increasing peak has been reported previously [39] and represents when the Ni oxides transitions from no spontaneous charge transfer with positive to spontaneous charge transfer case with negative . This indicates that ligand holes are already present and stable in pristine LiNiO2 (Ni3+) and that compounds with increasing formal oxidation state (Li2NiO3—Ni4+) relate to the presence of additional ligand holes. The invariant Ni 3d band intensity for all the Ni oxides in the RPES experiments further corroborates that increasing in formal oxidation states doesn’t affect the Ni 3d occupancy and rather links with the formation of ligand holes. This is why the O K-edge pre-edge peak increases with increasing oxidation states.

### v3

@@ Abstract
Describing Li-ion battery positive electrodes in terms of distinct transition metal or oxygen redox regimes can lead to confusion in understanding metal-ligand hybridisation, oxygen dimerisation, and degradation. There is a pressing need to study the electronic structure of these materials and determine the role each cation and anion plays in charge compensation. Here, we employ transition metal L-edge X-ray Resonance Photoemission Spectroscopy in conjunction with Single Impurity Anderson models, Self-consistent Real Space Multiple Scattering spectral simulations, and Dynamical Mean-Field theory calculations to directly evaluate the redox mechanisms in (de-)lithiated battery electrodes. This approach reconciles the redox description of two canonical positive electrodes—LiMn0.6Fe0.4PO4 and LiNiO2—in terms of varying degrees of charge transfer using the established Zaanen-Sawatzky-Allen framework, common to condensed matter physics. In LiMn0.6Fe0.4PO4, the absence of charge transfer means capacity arises due to the depopulation of metal 3d states, i.e. conventional metal redox. Whereas, in LiNiO2, charge transfer dominates and redox occurs through the formation and elimination of ligand hole states. This work clarifies the role of oxygen in Ni-rich system and provides a framework to explain how capacity can be extracted from oxygen-dominated states in highly covalent systems without needing to invoke dimerisation.
@@
Li-ion battery positive electrodes, Electronic structure, X-ray spectroscopy, X-ray simulationsOxygen loss limits the performance of both Li-excess and stoichiometric Ni-rich positive electrodes. [1, 2, 3] Efforts to suppress it are often conflated with oxygen redox, which is considered to involve O2 formation in Li-excess compounds. [4] This has prompted the field to re-examine oxygen participation in redox more broadly, leading to two conceptually distinct formalisms over the last decade: non-bonding O 2p orbitals and metal-ligand orbital hybridisation. [5] The former has been widely used to explain O2 formation observed in RIXS experiments, [4] yet compounds that challenge this interpretation also show oxygen-dimer features at the top of charge. [6, 7, 8, 9] Recent studies further question whether oxygen-dimer RIXS signals correlate with reversible capacity [10] while others propose alternative structural pathways to the O2 formation as the dominant factor of the electrochemistry of Li-excess layered materials. [11, 12] Within this conflicting view, oxygen participation via hybridisation is increasingly supported, with beyond density function theory studies confirming its role to deliver capacity from the onset of delithiation, [13] validated by in operando X-ray measurements [14]. As a result, directly probing oxygen participation through orbital rehybridisation can thus inform computational models and reconcile conventional redox frameworks with this emerging hybridisation perspective, advancing our understanding of redox mechanisms.
Equation 1 illustrates the end points of LiNiO2 delithiation, where charge compensation raises the Ni oxidation state. Linking this evolution to the underlying electronic structure is essential to understand oxygen loss. Generally, the classical redox model ties oxidation states to changes in transition metal (TM) 3d charge density, assuming a fully ionic character of the TM-O bonds. [15, 16] This framework views oxidation as discrete changes in TM 3d orbital occupancy. [17] LiMn0.6Fe0.4PO4 is a canonical example of such ionic redox.
Charge compensation in Equation 2 produces a Fe2+/Fe3+ redox pair that directly reflects variations in Fe 3d occupancy. By contrast, growing experimental evidence shows that this ionic picture fails to describe LiNiO2 because of its more complex electronic structure shaped by bond-disproportionation, [18] intersite charge fluctuations, [19] and strong metal-ligand hybridisation. [20] Under these influences, LiNiO2 redox pathways deviate from the simple ionic view. Theoretical work has begun to address such complexities [21, 12, 22], but further experimental validation is needed to establish a general framework that captures electronic intricacies across cathode materials and reconciles emerging and conventional redox views, with particular attention to oxygen’s role in charge compensation.
Most studies on Ni-rich layered compounds to date infer the redox mechanism indirectly, i.e. by probing core-level or unoccupied orbitals that are not themselves directly involve in charge compensation. In the classical ionic redox view, this indirect approach is generally sufficient. However, in compounds such as LiNiO2, where the electronic structure is more complex, such indirect interpretations may fail, giving room to misleading conclusions. For instance, during Li deintercalation of LiNiO2 positive electrodes, hard X-ray photoelectron spectroscopy (HAXPES) shows that Ni 2p peaks narrow but do not shift in energy as would be expected for ionic compounds. [23] In Ni K-edge X-ray absorption near-edge spectroscopy (XANES), the shift to higher energies of the raising edge—typically associated with Ni oxidation—halts despite continued capacity extraction, inconsistent with a model based solely on Ni oxidation. [24, 25, 8] Raman spectroscopy likewise reveals evolving E and A1g peak intensities that suggest a continuous change in Ni-O bond covalency rather than the spectral behaviour characteristic of ionic bonds. [19] Finally, Ni L-edge resonant inelastic X-ray scattering (RIXS) indicates a dominant Ni 3d8 character, rather than the expected 3d7 configuration implied by classical ionic redox. [19] Taken together, these results highlight the need for methods that circumvent the limitations of indirect electronic probes and instead directly resolve the character and evolution of redox-active orbitals. Such an approach could advance our understanding of redox mechanisms, consistent with the emphasis in recent computational studies. [26] Motivated by this need, our work implements a methodology that directly interrogates the orbitals involved in charge compensation. By unambiguously identifying their elemental and orbital character, this method avoids the concerns of indirect probes and provides a clearer understanding of the redox mechanism in electronically complex materials.
This approach is TM L-edge X-ray Absorption Spectroscopy (XAS)-assisted Resonant Photoemission Spectroscopy (RPES). [27, 28] By performing valence-band photoemission spectroscopy with X-ray energies tuned to TM L-edge transitions, RPES directly reveals which elements participate in redox and—when combined with the correct electronic assignment—how their electronic structure evolves. This assignment is provided by Single Impurity Anderson (SIA) models [29], which determine whether the selected transitions involve simple t2g/eg states or more complex features such as intersite charge fluctuations, bond-disproportionation, or metal-ligand hybridisation, thereby ensuring that resonant enhancement amplifies the relevant orbital character in the spectra. This reciprocal experiment-theory synergy enables direct probing of how intricate electronic structures dictate the redox mechanism, providing a material-agnostic framework for tracking orbital-level redox processes across diverse compounds and states of charge (SoC).
@@
Here, we used the TM L-edge XAS-assisted RPES approach to directly study charge compensation mechanisms in the context of the simple ionic picture and more complex electronic structures. We used this approach on two canonical systems, LiMn0.6Fe0.4PO4 (LMFP64) and LiNiO2 (LNO). The high quality and electrochemical performance of the LMFP64 material and electrodes used in this work have been proven previously. [30] The LNO electrodes of this work exhibit the expected accelerated electrochemical performance decay [25] as a consequence of cycling above its O-loss cell potential threshold at 4.1 V vs. LiLi+, which we used to intentionally condition the material with relevant voltage windows for this work. LMFP64 displays the classical redox picture due to an intense induction effect from the P ions. [31] In contrast, by carefully considering bond-disproportionate site, intersite charge fluctuation, and metal-ligand orbital hybridisation effects, a different redox view emerges in LNO more linked to a local site and ligand hole reconfiguration. Our findings provide critical insights to the electronic structure and a pathway to understanding oxygen redox.
RPES: Classical and Non-classical Redox
The Classical Redox in LMFP64
@@
The Fe redox mechanism in LMFP64 represents the archetypal conventional model, where delithiation corresponds to electron depopulation from Fe 3d orbitals. Within our approach, we use this material as a reference to establish the spectral fingerprints of the conventional redox picture. To this end, we prepared a set of LMFP64 half-coin cell formats to charge the cells to targeted cell potentials as indicated by the coloured circles in Figure 1a. In general, LMFP64 undergoes unexpected structural transformations under fast charge rates, [32] which can affect intercalation reactions and, consequently, the rate of Fe-ion charge compensation. [33] To minimise these effects, we employed a slow C/20 charging rate (7.93 mA/g), as described in the Methods section. This guarantees a cell potential profile with negligible polarisation and, as a result, a better approximation to the open circuit voltage profile (the true cell potential profile of cathode materials). [3]
@@
The top panel in Figure 1b shows the experimental Fe L-edge spectrum of the pristine electrode (no cycling and no exposure to electrolyte) and the SIA simulated Fe L-edge spectrum. We performed the Fe L-edge simulations without explicit Fe 3d - O 2p ligand hybridisation (to guarantee an ionic picture) and with six Fe 3d electrons (to generate the expected 2+ valence). The remarkable match between the Fe L-edge experimental and simulated spectra confirms the ionic picture best describes the LMFP64 electronic structure; this means that an increase in the oxidation state inherently implies a reduction in the electron occupancy of the Fe 3d orbital. To simulate the fully charged condition at 4.5 V, we deliberately reduced the number of Fe 3d electrons from six to five in the simulation while keeping the no Fe 3d - O 2p ligand hybridisation condition. Spectral comparison shows again a remarkable match of the simulated and experimental Fe L-edge spectra (bottom panel in Figure 1b). Combining these findings for the pristine and 4.5 V electrodes strongly suggests that the Fe redox mechanism involves a reduction in the number of Fe 3d electrons, leading to an increase in the Fe oxidation state from 2+ to 3+ in LMFP64 during charge. Notice that in the 704 - 709 eV energy range, Fe 3d-orbitals are the dominant feature of the Fe L-edge XAS transitions during charge (Figure 1c). Resonating in this range with RPES experiments will reveal the LMFP64 redox mechanism as a depopulation of the Fe 3d orbitals while the cell charges, proving the classical redox view in LMFP64.
@@
Figure 1d shows RPES results at progressively increasing cell potential (pristine, 3.41, 3.80, 4.05, and 4.50 V, respectively) in the Fe 3d dominated energy range. With increasing potentials, we observe the RPES Fe 3d intensity at 2 eV binding energy decreases after the Fe2+/3+ potential plateaus in Figure 1a, experimentally confirming the conventional redox picture and that Fe oxidation state changes in LMFP64 links to the valence-counting formalism on the Fe 3d orbital. A similar procedure was employed to investigate the participation of the Mn sites in the redox mechanism. The analysis prompted further investigation of a possible degradation mechanism stemming from the Mn2+/3+ redox couple that we will explore in a subsequent more in-depth manuscript regarding LMFP64 degradation.
LNO: Classical Redox?
If Ni 3d valence-counting governs the redox mechanism in LNO cells, Ni L-edge RPES experiments should reveal spectral evolution similar to that observed for Fe. To test this hypothesis, we assembled half coin cells (vs. LiLi+) and charged them at C/20 (C=220 mA/g as described in the Methods section) to 4.2 V and later discharged them at 3.0 V, Fig. 2a.
Contrary to the Fe case in LMFP64, Ni L-edge simulations in the ionic picture (i.e., without explicit Ni 3d-O 2p hybridisation) show poor agreement with experiment. Both the pristine state, simulated with 7 electrons in the Ni 3d orbital, and the charged state, simulated with 6 electrons, exhibit significant mismatch with bulk-sensitive Ni L-edge spectra in TFY mode (Fig. 2b). This indicates that more complex electronic structures govern LNO redox during cycling, with electron depopulation arising from states beyond the simple Ni 3d orbital.
This more complex redox evolution is revealed through a strategic integration of RPES and theoretical frameworks, representing—to our knowledge—the first direct evidence of metal-ligand redox. In Fig. 2c, we present valence-band XPS spectra of LNO electrodes at 4.2 V and after discharge to 3.0 V. While changes in the topmost valence band clearly track the electrochemical state, these changes alone lack both element specificity and orbital resolution. To resolve the orbital character, we employed DMFT and GW0 calculations (Fig. 2d) alongside Ni L-edge and O K-edge RPES measurements (Fig. 2e). Despite their distinct theoretical frameworks, both DMFT and GW0 consistently assign the valence-band edge primarily to a mixed Ni 3d-O 2p character, with the Ni 3d states positioned deeper, around 10 eV. This assignment suggests that LNO redox involves spectral evolution within this strongly hybridised Ni 3d-O 2p orbital.
RPES measurements directly confirm the theoretical assignment of Ni 3d-O 2p hybridisation and substantiate its role in governing LNO redox. Ni L-edge RPES shows an intense, electrochemically invariant feature at 10 eV, consistent with theory predicting deeply bound Ni 3d states. These observations strongly indicate that the dominant Ni 3d electron occupancy remains unchanged during cycling, i.e., throughout LNO redox. When resonating at the O K-edge pre-edge, the RPES spectrum reveals a strong feature at the same 10 eV, confirming substantial Ni 3d-O 2p hybridisation. Crucially, only the O K-edge RPES exhibits clear discharge-induced changes, particularly in the top valence states. Building on the Fe case, where such changes signify electron occupancy variation, the combined Ni L-edge and O K-edge RPES trends indicate that LNO redox fundamentally involves electron population changes on ligand states in strongly hybridised orbitals.
Collectively, our results reveal an intriguing shift in the redox mechanism, from metal-centred in LMFP64 to ligand-dominated through cooperative Ni-O hybridisation, and thereby open new avenues of inquiry: why does this shift occur, what drives it, and how does it shape our broader understanding of redox chemistry?
Electronic Structures: From Metal-Centred to Hybridised
The shift in the redox mechanism described in the previous section arises from electronic structures, which evolve toward greater complexity with increasing formal oxidation state of the metals. At higher oxidation states, the stronger electron affinities of the metals and the weaker electronegativities of the ligands combine to produce compounds with more intricate electronic structures, [16] thereby defining the pathways through which redox occurs. Using our approach, we designed experiments to identify the spectroscopic signatures of this electronic evolution and lay the foundation to bridge fundamental insights with cell-level redox performance, establishing a reliable from-quantum-to-cell loop for a comprehensive understanding of LNO redox in the next section.
@@
To this purpose, we prepared a set of four Ni-oxide compounds (LaNiO2, NiO, LNO and Li2NiO3), where the Ni ion is oxidised in different formal oxidation states (i.e. Ni1+, Ni2+, Ni3+ and Ni4+). The Ni-O ligand hybridisation is minimal in LaNiO2 [34] while playing a substantial role in NiO. [35] LNO is better described with intersite charge fluctuations, explicit TM 3d - O 2p ligand hybridisation, and bond-disproportionate NiO6 sites. [23] Li2NiO3 is the endpoint of Li-excess and its electronic structure is better described by intersite charge fluctuations and explicit TM 3d - O 2p ligand hybridisation. [36] Therefore, our approach applied to the Ni oxide material set will show the spectral difference of the transition towards more complex electronic structures.
Figure 3a shows Ni L-edge XAS spectra of as-synthesised LaNiO2, NiO, LNO, and Li2NiO3 ,with the corresponding RPES measurements in Figure 3b. Given the shallow probing depth of RPES measurements, Ni oxides prone to surface reconstruction can produce spectral convolution between surface species and the subsurface region. Figure S1 identifies that surface spectral contribution concentrates at 853 eV and energies 854 eV are redox-related originating from subsurface regions beneath the surface-reconstructed layer. We thus focused our RPES experiments and analysis on Ni L-edge transitions with energies 854 eV.
Building on the LMFP64 case, where the redox-related spectral signature appears at the topmost rising edge, our analysis shows that the Ni 3d bands (indicated by arrows in Figure 3b) are not redox-active, as they lie deeper in energy (8-10 eV) across the Ni-based oxide set, consistent with the LNO cathode observations discussed in the previous section. Instead, the states located near EF—within the redox-active regions, highlighted by the white square in Figure 3b—are redox-relevant. Resolving the nature of these states, by identifying XAS transitions through which they resonate in Ni L-edge RPES (855-865 eV), is key to understanding the evolution from metal-centred to hybridised electronic structures and the associated redox processes in Ni compounds. By combining dynamical mean-field theory (DMFT) for orbital occupancy analysis, Ni L-edge transition simulations within SIA models, and O K-edge calculations using real-space multiple-scattering theory in an interacting Green’s function basis (FEFF), we provide strong multi-level evidence that spontaneous ligand-to-metal charge transfer drives the electronic structure evolution, thereby defining the redox-active bands that ultimately dictate the redox pathways during Ni-based cell operation.
As the formal oxidation state increases, the cooperative effect of stronger electron affinity and weaker electronegativity reaches a threshold at which, in strongly hybridised orbitals, charge redistributes between Ni and O, allowing partial charge transfer—also referred to as charge fluctuation—to stabilise. Experimentally, this phenomenon is parametrised by a factor , with positive values indicating the stage prior to, and negative values marking the onset of, spontaneous ligand-to-metal charge transfer. As demonstrated by SIA-based Ni L-edge simulations of the positive- NiO system (Fig. S2), these charge fluctuations dominate transitions in the Ni L-edge XAS energy range (855-865 eV), where resonance in RPES gives rise to the Ni-O hybridised orbitals at the top of the VB, i.e. the redox-relevant states of Ni compounds. Explicit inclusion of Ni 3d-O 2p ligand hybridisation in the simulations (Fig. S2b) was critical to resolving the features within this 855-865 eV range. However, because remains positive for NiO, this fluctuation is not spontaneous but instead arises from X-ray excitation—commonly referred to as the Zhang-Rice bound state [35]—and is therefore not intrinsic to NiO’s electronic structure, explaining why NiO lacks spectral features of charge-transfer-related states in the Ni L-edge RPES resonance between 855-865 eV. We therefore label the 855-865 eV window as the -dominant region, since its RPES resonance links directly to hybridised orbitals with spontaneous charge-transfer character, which are expected to be prominent in negative- systems.
Indeed, the pronounced Ni L-edge RPES features resonating in this -dominant range for LNO and Li2NiO3 indicate that the transition to the negative- regime occurs between Ni2+ and Ni3+. Thin films prepared to simulate this Ni oxidation progression further support this view, implying expectations of non-conventional redox in Ni3+ systems arising from their negative- character. DMFT orbital occupancy analysis (Fig. S3) confirms this spontaneous charge migration beyond Ni2+, stabilised through the gradual formation of electronic “holes” or in the O 2p ligand state within the hybridised orbitals. Spectral measurements (Fig. S3) validate these theoretical predictions, revealing hole formation as intensity migration from the topmost occupied states (hybridised Ni 3d-O 2p ligand) to the lowest unoccupied states (O K-edge pre-peak). The rising O K-edge pre-peak thus serves as a clear manifestation of ligand hole formation and stabilisation in the hybridised Ni 3d-O 2p ligand orbital, consistent with previous reports. [37, 38]
Following this interpretation for Fig. 3b-d, LNO and Li2NiO3 exhibit clear signatures of negative- systems. Their formal oxidation states exceed Ni2+, placing them well beyond the positive-to-negative transition threshold. This results in stabilised ligand holes, as confirmed by DMFT calculations in both compounds. Experimentally, the presence of ligand holes manifests as pronounced Ni L-edge intensity in the -dominated range. The progressive growth of the O K-edge pre-edge feature also reflects the DMFT-predicted increase in ligand-hole numbers as oxidation proceeds from Ni3+ to Ni4+, consistent with the corresponding growth of the pre-edge peak in FEFF-based simulations. Together, these observations establish the O K-edge pre-edge as a robust descriptor of ligand holes, endorsing ligand-hole formation as the mechanism driving evolution toward higher Ni oxidation states. Crucially, the strong consistency in describing ligand-hole formation across three distinct theoretical frameworks—DMFT, multiplet effects in SIA, and real-space correlation in GW0 (FEFF)—reconciles these approaches and underscores both the robustness of this description and the central role of ligand holes in dictating the redox pathways of high Ni oxidation-state systems.
This integrated experimental and theoretical analysis firmly establishes LNO and Li2NiO3 as prototypical negative- systems, in which ligand-hole is already present in the pristine state and its formation and annihilation is the dominant mechanism governing the evolution across the full range.